OpenStudy (anonymous):

What is the probability that n tails in a row with a coin for which the probability of heads is p? Given a fraction , with 0 < alpha < 1, show that the number of flips one must perform to be  sure of obtaining at least one head is n =ln(1-alpha)/ln(1-p)  .

OpenStudy (anonymous):

answer to first question is that if the probability of heads is \(p\) the then probability of tails is \(1-p\) and since the events are presumed to be independent, the probability of n tails in a row is \((1-p)^n\)

OpenStudy (anonymous):

i don't understand what the second part of the question says. looks like something is missing

OpenStudy (anonymous):

given a fraction alpha , with 0 <alpha<1, show that the number of flips one must perform to be alpha sure of obtaining atleast one head is n=ln(1-alpha)/ln(1-p)

OpenStudy (anonymous):

ok i am thinking out loud, so maybe this is wrong. probabiliy you get at least one head in \(n\) tosses is one minus probability you get all tails is \(1-(1-p)^n\)

OpenStudy (anonymous):

if you put \(X_n\) as the number of heads in n tosses, then \(P(X_n\geq 1)=1-(1-p)^n\) and you want to solve \[P(X\geq 1)\geq \alpha\] ( i am not sure why alpha has to be a fraction)

OpenStudy (anonymous):

so i think this comes down to solving \[1-(1-p)^n \geq \alpha\]

OpenStudy (anonymous):

oh and it works exactly. not bad for first thing in the morning, huh?

OpenStudy (anonymous):

first off did what i write make sense? and secondly can you solve this for \(n\)?

OpenStudy (anonymous):

how abou introducing ln both sides

OpenStudy (anonymous):

yes the solution starts with \[(1-p)^n\leq 1-\alpha \] then the change of base formula gets it in one step

OpenStudy (anonymous):

still not sure why \(\alpha\) has to be a "fraction"

OpenStudy (anonymous):

maybe ase 's a probability and it is between 0 and 1

OpenStudy (anonymous):

yes but no one says it has to be a "fraction" it could be \[\sqrt{.5}\] for example. there are an infinite number of numbers between 0 and 1 that are not "fractions" no matter, i think this is done right?

OpenStudy (anonymous):

ya get you point

OpenStudy (anonymous):

oh it occurs to me that there is one more thing to say, although it may be obvious \[(1-p)^n\leq 1-\alpha\] \[n\ln(1-p)\leq \ln(1-\alpha)\]change of base gives \[(1-p)^n\leq 1-\alpha \iff n \[n\geq \frac{\ln(1-\alpha)}{\ln(1-p)}\] switch the inequality because evidently \(\ln(1-p)\leq 0\)

OpenStudy (anonymous):

\[(1-p)^n\leq 1-\alpha \iff n\geq \frac{\ln(1-\alpha)}{\ln(1-p)}\] is what i meant

OpenStudy (anonymous):

will i be wromg if i put \[\le\]

OpenStudy (anonymous):

hell yes

OpenStudy (anonymous):

y

OpenStudy (anonymous):

forget about the math for a second, think about what it is asking how many tosses before you are more than \(\alpha\) certain of getting at least one head. clearly your inequality should look line \(n\geq \text{something}\)

OpenStudy (anonymous):

in other words n has to be "at least" some number, not "less than" some number. that wouldn't make any sense

OpenStudy (anonymous):

oh i get what u mean now

OpenStudy (anonymous):

we start with \[(1-p)^n\leq 1-\alpha\] and since the log is a one to one and increasing function the inequality is preserved when we take the log and get \[n\ln(1-p)\leq \ln(1-\alpha)\]

OpenStudy (anonymous):

but now to solve for \(n\) we must divide both sides by \(\ln(1-p)\) and since \(1-p<1\) we know \(\ln(1-p)<0\) i.e. it is negative. when you divide by a negative number you have to change the sense of the inequality so you get \[n\geq \frac{\ln(1-\alpha)}{\ln(1-p)}\] good thing too, otherwise the answer wouldn't make any sense at all